if difference between x and y is 12 and their ratio is 3:7. what is the value of y?
Answer by richwmiller(17219)

x-y=12
x=12+y
x/y=3/7
x=3y/7
x=12+y
3y/7=12+y
3y=84+7y
-4y=84
y=-21
x=-9
or
y-x=12
y=12+x
x/y=3/7
y=7x/3
12+x=7x/3
36+3x=7x
36=4x
9=x
y=21